Let Pun be a series of arbitrary terms. For all n ∈ N, let pn =

1

2

(un + |un|) and qn =

1

2

(un − |un|). Show

that

(a) If Pun is absolutely convergent, then both Ppn and Pqn are convergent.

(b) If Pun is conditionally convergent, then both Ppn and Pqn are divergent.


1
Expert's answer
2021-05-21T03:44:46-0400

"P_{u_n} = \sum_{n=1}^{\infty} u_n \\
p_n = u_n + |u_n| \\
q_n = u_n - |u_n|"

a) Suppose "P_{u_n} = \sum_{n=1}^{\infty} u_n \\" is absolutely convergent. This implies that "\sum_{n=1}^{\infty} |u_n|" is convergent and hence "\sum_{n=1}^{\infty} u_n" is also convergent (by the convergence of "\sum_{n=1}^{\infty} |u_n|" )

Consider,

"P_{p_n}= \sum_{n=1}^{\infty} p_n = \sum_{n=1}^{\infty} (u_n + |u_n| )= \sum_{n=1}^{\infty} u_n + \sum_{n=1}^{\infty} |u_n|"

And since the sum of two convergent series is convergent, we can safely conclude that "P_{p_n}" is convergent as required.

Also, consider

"P_{q_n} = \sum_{n=1}^{\infty} q_n = \sum_{n=1}^{\infty} (u_n - |u_n| )= \sum_{n=1}^{\infty} u_n -\sum_{n=1}^{\infty} |u_n|"

And since the difference of two convergent series is convergent, we can safely conclude that "P_{q_n}" is convergent as required.


b) Suppose "P_{u_n} = \sum_{n=1}^{\infty} u_n \\" is conditionally convergent. This implies that "\sum_{n=1}^{\infty} |u_n|" is divergent and "\sum_{n=1}^{\infty} u_n" is convergent.

Consider,

"P_{p_n} = \sum_{n=1}^{\infty} p_n = \sum_{n=1}^{\infty} (u_n + |u_n| )= \sum_{n=1}^{\infty} u_n + \sum_{n=1}^{\infty} |u_n|"

Since, the sum of a convergent series and a divergent series yields a divergent series, we can safely conclude that "P_{p_n}" is a divergent series as required.

Also, consider

"P_{q_n} = \sum_{n=1}^{\infty} q_n = \sum_{n=1}^{\infty} (u_n -|u_n| )= \sum_{n=1}^{\infty} u_n -\sum_{n=1}^{\infty} |u_n|"

Since, the difference between a convergent series and a divergent series yields a divergent series, we can safely conclude that "P_{q_n}" is a divergent series as required.